4) You performed matings with other flies from question 3 and found some unusual results. The 1600 F2 progeny are the result of crossing the Fl progeny to each other. P1: thick wings, white bristles, long antenna female x skinny wings, red bristles, aba male Fl: all thick wings, red bristles, long antenna flies F2: 910 thick wings, red bristles, long antenna flies 290 skinny wings, red bristles, aba flies 110 skinny wings, white bristles, aba flies 280 thick wings, white bristles, long antenna flies 3 thick wings, red bristles, long antenna flies 4 skinny wings, red bristles, aba flies 1 thick wings, white bristles, long antenna flies 2 skinny wings, white bristles, aba flies A. What is the inheritance pattern of the traits? B. Which, if any, of the genes are linked? What are the map distances between the linked genes?Only questoin 4 but it refers to question 3 which is why i put question 3 as well.
Answer #1

B. There is incomplete linkage has occurred.

910 thick wings (T), red bristles (R), long antenna flies (L) – T R L (wild type)

110 skinny wings (t), white bristles (r), aba flies (l) – t r l (wild type)

290 skinny wings, red bristles, aba flies – t R l (double cross over [D.C.O])

280 thick wings, white bristles, long antenna flies – T r L (double cross over [D.C.O])

4 skinny wings, red bristles, long antenna flies – t R L (single cross over [S.C.O] I)

2 skinny wings, white bristles, long antenna flies – t r L (single cross over [S.C.O] I)

3 thick wings, red bristles, aba flies – T R l (single cross over [S.C.O] II)

1 thick wings, white bristles, aba flies – T r l (single cross over [S.C.O] II)

Total progeny – [ 910 + 110 + 290 + 280 + 4 + 2 + 3 + 1] = 1600

Frequency of double cross over – [(570/1600)*100] = 35.625

Frequency of S.C.O I – [(6/1600)*100] = 0.375

Frequency of S.C.O II – [(4/1600)*100] = 0.25

Map distance between T and R is – S.C.O I + D.C.O = 35.625 + 0.375 = 36 m.u

Map distance between R and L – S.C.O II + D.C.O = 35.625 + 0.25 = 35.875 m.u

T---------------36 m.u-----------R-----------35.875 m.u-----------L